# ECLARE: Efficient cross-planar learning for anisotropic resolution enhancement

### v2

@@ Abstract
In clinical imaging, magnetic resonance (MR) image volumes are often acquired as stacks of 2D slices with decreased scan times, improved signal-to-noise ratio, and image contrasts unique to 2D MR pulse sequences. While this is sufficient for clinical evaluation, automated algorithms designed for 3D analysis perform poorly on multi-slice 2D MR volumes, especially those with thick slices and gaps between slices. Super-resolution (SR) methods aim to address this problem, but previous methods do not address all of the following: slice profile shape estimation, slice gap, domain shift, and non-integer or arbitrary upsampling factors. In this paper, we propose ECLARE (Efficient Cross-planar Learning for Anisotropic Resolution Enhancement), a self-SR method that addresses each of these factors. ECLARE uses a slice profile estimated from the multi-slice 2D MR volume, trains a network to learn the mapping from low-resolution to high-resolution in-plane patches from the same volume, and performs SR with anti-aliasing. We compared ECLARE to cubic B-spline interpolation, SMORE, and other contemporary SR methods. We used realistic and representative simulations so that quantitative performance against ground truth can be computed, and ECLARE outperformed all other methods in both signal recovery and downstream tasks. Importantly, as ECLARE does not use external training data it cannot suffer from domain shift between training and testing. Our code is open-source and available at https://www.github.com/sremedios/eclare.
@@ 1 Introduction
To address these concerns, we propose ECLARE (Efficient Cross-planar Learning for Anisotropic Resolution Enhancement), a deep learning algorithm to super-resolve multi-slice 2D MR volumes along the through-plane direction. ECLARE builds on a succession of self-SR approaches [33, 34, 35, 36], in which all training data is extracted from the HR in-plane slices within a LR image volume. However, ECLARE adds several crucial features to properly model slice acquisition, account for non-integer super-resolution factors, and incorporate efficient network design elements. In particular, we use ESPRESO [30], a GAN-based algorithm previously developed by our team, to estimate the relative slice selection profile. We use this profile to model slice thickness and gap and to simulate training data from HR in-plane patches extracted from the anisotropic input volume. These data are used to train an efficient SR neural network architecture based on “wide activation SR” (WDSR) [37]. The model is then applied to both canonical through-planes, yielding two super-resolved volumes that are then averaged to create the final output. Underlying all interpolation operations in our approach is a novel field-of-view (FOV)-aware interpolation algorithm that respects the image coordinate system and target sample spacing. The main contributions of this paper are the following:
  • We propose an FOV-aware interpolation to resample an image to a new grid while preserving a desired sample spacing and maintaining the location of the FOV.
@@ 2 Related Work
Figure 2: A 1D demonstration of the differences between interpolation implementations. The input has 6 pixels and its FOV center is located at 2.5, indicated by the red dashed line. Its sample spacing is 1, indicated by the purple arrow. Its extents occur at -0.5 and 5.5, indicated by the blue dashed line. In this demonstration, the desired sample spacing for upsampling is 0.7, and the desired sample spacing for downsampling is 1.429. The implementation scipy.zoom [38] maintains the FOV center and boundary sample locations, but changes the sample spacing to 0.625 and 1.667 for upsampling and downsampling, respectively. The implementation torch.nn.functional.interpolate [39] anchors the FOV extent to start at 0.5, shifting the center from the desired FOV center. Our proposed method (Sec. 3.3) maintains both the sample spacing and the FOV center.
@@
Implicit neural representations (INR) [43, 44, 45] learn the mapping from coordinate to intensity. Since coordinates can be sampled at any resolution, this has inspired several works related to arbitrary-scale SR [46, 45, 47]. Recently, multi-contrast INR [48] was proposed to super-resolve subjects with multiple available contrasts from the same session, ideally acquired at different orientations (e.g., sagittal $T_{1}$-w and coronal T2-w). However, that method also supports the use of a single contrast, training in a similar manner to self-SR methods.
@@ 3 Methods
Our proposed ECLARE improves resolution along the through-plane axis of an anisotropic 2D-acquired multi-slice MR image volume. To explain ECLARE, we first model the relationship between HR and LR volumes (Sec. 3.1). We then describe how to approximate the slice selection profile (Sec. 3.2) and FOV-aware interpolation (Sec. 3.3). We approximate the slice selection profile with ESPRESO [30] and use it to train a CNN with in-plane HR patches and their simulated LR counterparts (Sec. 3.4). Finally, we show how ECLARE uses this 2D CNN to super-resolve the 3D volume (Sec. 3.5) and describe implementation details (Sec. 3.6).
@@ 3.2 Approximating the imaging process
3.3 FOV-aware interpolation
While interpolation is routine in digital signal processing, the exact implementation has several freedoms. ECLARE requires both downsampling (for training data simulation) and upsampling (for inference) and we would like both to be stable with respect to the FOV. We consider three conditions that are desirable to meet with respect to the FOV of the signal:
@@ 3.3 FOV-aware interpolation
Interpolation algorithms cannot meet all of these conditions in general since only a discrete number of samples is permitted. However, since it is possible to satisfy two of these conditions, a compromise on the third is necessary. We quickly review popular modern implementations. The method scipy.ndimage.zoom [38] meets conditions 1) and 3). It adjusts the sample spacing to preserve the FOV center while locking the locations of boundary samples. The method torch.nn.functional.interpolate [39] meets only condition 2). It preserves the sample spacing, but neither the FOV center nor the boundary sample locations. Instead, it anchors the extent on the first sample. We place more importance on the center of the FOV and the sample spacing and therefore desire conditions 1) and 2). We summarize the interpolation method proposed in [52] that maintains the FOV center and sample spacing while sacrificing the precise location of the boundary samples. A comparison of the three implementations is shown in Fig. 2.
@@
As in [52], we introduce a fixed coordinate system for the signal before and after interpolation. Let the position of the first pixel be centered at zero, let $N$ be the number of pixels before interpolation, and let the spacing before interpolation be one. The width of each pixel is equal to the spacing; before interpolation it is therefore also one. The first pixel $p_{0}$ is centered at zero and the last pixel $p_{N-1}$ is centered at $N-1$. Since each pixel has a width of one, $p_{0}$ extends from $-0.5$ to $0.5$ and $p_{N-1}$ extends from $N-1.5$ to $N-0.5$. Therefore the entire image extends from $-0.5$ to $N-0.5$.
[…]
@@
Accordingly, we notate the first pixel after interpolation as $p^{\prime}_{0}$ and the last pixel as $p^{\prime}_{N^{\prime}-1}$.
In order to center around the same FOV, the center position of the entire image should be identical before and after interpolation. Since the center of the FOV is the halfway point between the first and last pixels, we can write
@@
Finally, we can calculate the new extents of the entire image after interpolation. The first pixel extends from $p^{\prime}_{0}-\frac{d}{2}$ to $p^{\prime}_{0}+\frac{d}{2}$ and the last pixel extends from $p^{\prime}_{N-1}-\frac{d}{2}$ to $p^{\prime}_{N-1}+\frac{d}{2}$. Thus, the image extents are $p^{\prime}_{0}-\frac{d}{2}$ to $p^{\prime}_{N-1}+\frac{d}{2}$.
The result of this method yields an FOV-aware interpolation implementation that maintains the FOV center and guarantees the desired pixel spacing. This is true in both upsampling and downsampling scenarios as demonstrated in Fig. 2.
@@ 3.4 Training a deep learning model with paired data
We now develop a relationship between in-plane and through-plane data and train a SR network. Paired LR and HR data are necessary for supervised learning. We extract HR 2D patches from in-plane and produce blurred 2D patches using Eq. 1 with the learned $\tilde{h}$ from ESPRESO. These blurred 2D patches are then downsampled using the FOV-aware interpolation described in Sec. 3.3. Thus, each LR patch has the appropriate simulated slice thickness, separation, and FOV.
@@
[…]
@@ 4.1 Data and evaluation metrics
Our experiments began with HR isotropic 3D-acquired image volumes of resolution 1×1×1 $1\times 1\times 1\text{~{}mm}^{3}$. These images were then degraded with Eq. 1, setting $h$ to the shape of the slice excitation profile created by the Shinnar-Le Roux (SLR) pulse design algorithm [50, 51], $s$ to the desired slice spacing (described below), and $x$ to a 3D-acquired isotropic volume. The implementation of the SLR algorithm we use is from the sigpy library [55].
[…]
@@ 5 Discussion and conclusions
In this paper we presented ECLARE, a novel self-SR algorithm for anisotropic volumes that trains solely on data extracted from in-plane slices and thus requires no external training data. We presented a model of the imaging process in 2D multi-slice MRI and used ESPRESO [30] to estimate the slice selection profile. With an estimated slice profile, slice thickness and slice separation are both modeled, allowing for training data simulation from the HR in-plane data already present in the anisotropic volume. ECLARE uses a neural network architecture based on WDSR to efficiently train weights from scratch in a short amount of time, taking approximately 5 minutes on a single NVIDIA V100 GPU with a memory footprint under 2 GB. The result is a SR technique that respects the LR image contrast and presence of pathology, does not require external training data, and produces HR image volumes with the correct digital resolution and FOV. When compared to four other approaches on two simulated datasets, our approach achieved superior quantitative and qualitative results with respect to signal recovery metrics PSNR and SSIM as well as a downstream task metric—cDSC for brain region and MS lesion segmentation. This is true for all investigated resolutions except for $3\|0$, where SMORE performs best. This is likely due to two reasons: 1) SMORE does not have a model of slice gap, and therefore will perform best when no gap is present in the data; and 2) SMORE relies on pre-trained weights and was primarily tested and evaluated using data with resolutions $2\|0$ and $3\|0$.
We note that computation time is a primary limitation of ECLARE. ECLARE must fit a model for each multi-slice 2D MR volume. Although its training time is relatively fast compared to other self-SR methods, it is slower than pre-trained models that perform inference only. It would be possible to pre-train-then-fine-tune ECLARE as SMORE, but this introduces a tradeoff between time and relying on pre-training as a prior more than the in-plane data itself. Future work will investigate these tradeoffs in detail.

### v3

@@ Abstract
Purpose: In clinical imaging, magnetic resonance (MR) image volumes are often acquired as stacks of 2D slices with decreased scan times, improved signal-to-noise ratio, and image contrasts unique to 2D MR pulse sequences. While this is sufficient for clinical evaluation, automated algorithms designed for 3D analysis perform poorly on multi-slice 2D MR volumes, especially those with thick slices and gaps between slices. Super-resolution (SR) methods aim to address this problem, but previous methods do not address all of the following: slice profile shape estimation, slice gap, domain shift, and non-integer or arbitrary upsampling factors.
Approach: In this paper, we propose ECLARE (Efficient Cross-planar Learning for Anisotropic Resolution Enhancement), a self-SR method that addresses each of these factors. ECLARE uses a slice profile estimated from the multi-slice 2D MR volume, trains a network to learn the mapping from low-resolution to high-resolution in-plane patches from the same volume, performs SR with anti-aliasing, and respects the image FOV during resampling. We compared ECLARE to cubic B-spline interpolation, SMORE, and other contemporary SR methods. We used realistic and representative simulations on human head MR volumes so that quantitative performance against ground truth can be computed. Specifically, healthy $T_{1}$-w and people with MS $T_{2}$-w FLAIR datasets were used for evaluations. We used the peak signal to noise ratio and structural similarity index measure as signal recovery metrics. We additionally used two independent brain parcellation algorithms, SLANT and SynthSeg, to compute the consistency Dice similarity coefficient and the $R^{2}$ coefficient of determination, respectively, as comparison metics.
Results: For images with up to 5 mm of slice thickness and up to 1.5 mm of gap, ECLARE achieves greater mean PSNR and SSIM compared to other methods. In representative regions of interest, such as the ventricles, caudate, cerebral white matter, and cerebellar white matter, ECLARE performs comparably or better than other approaches. These trends are similar for both investigated datasets.
Conclusions: The use of slice profile estimation, FOV-aware resampling, and self-SR allowed ECLARE to robustly super-resolve anisotropic images without the need of external training data. Future work will investigate the utility of ECLARE on other organs, species, modalities, and resolutions. Our code is open-source and available at https://www.github.com/sremedios/eclare.
keywords:
inverse problem, magnetic resonance imaging, slice gap, super-resolution
[…]
@@ 1 Introduction
To address these concerns, we propose ECLARE (Efficient Cross-planar Learning for Anisotropic Resolution Enhancement), a deep learning algorithm to super-resolve multi-slice 2D MR volumes along the through-plane direction. ECLARE builds on a succession of self-SR approaches [19, 62, 38, 39], in which all training data is extracted from the HR in-plane slices within a LR image volume. However, ECLARE adds several crucial features to properly model slice acquisition, account for non-integer super-resolution factors, and incorporate efficient network design elements. In particular, we use ESPRESO [14], a GAN-based algorithm previously developed by our team, to estimate the relative slice selection profile. We use this profile to model slice thickness and gap and to simulate training data from HR in-plane patches extracted from the anisotropic input volume. These data are used to train an efficient SR neural network architecture based on “wide activation SR” (WDSR) [57]. The model is then applied to both canonical through-planes, yielding two super-resolved volumes that are then averaged to create the final output. Underlying all resampling operations in our approach is a novel field-of-view (FOV)-aware resampling algorithm that respects the image coordinate system and target sample spacing. The main contributions of this paper are the following:
  • We propose an FOV-aware resampling to regrid an image while preserving a desired sample spacing and maintaining the location of the FOV.
@@ 2 Related Work
Figure 2: A 1D demonstration of the differences between resampling implementations. The input has 6 pixels and its FOV center is located at 2.5, indicated by the red dashed line. Its sample spacing is 1, indicated by the purple arrow. Its extents occur at -0.5 and 5.5, indicated by the blue dashed line. In this demonstration, the desired sample spacing for upsampling is 0.7, and the desired sample spacing for downsampling is 1.429. The implementation scipy.ndimage.zoom [49] maintains the FOV center and boundary sample locations, but changes the sample spacing to 0.625 and 1.667 for upsampling and downsampling, respectively. The implementation torch.nn.functional.interpolate [36] anchors the FOV extent to start at 0.5, shifting the center from the desired FOV center. Our proposed method (Sec. 3.3) maintains both the sample spacing and the FOV center.
@@
Implicit neural representations (INR) [35, 42, 51] learn the mapping from coordinate to intensity. Since coordinates can be sampled at any resolution, this has inspired several works related to arbitrary-scale SR [46, 51, 27]. Recently, multi-contrast INR [34] was proposed to super-resolve subjects with multiple available contrasts from the same session, ideally acquired at different orientations (e.g., sagittal T1-w and coronal T2-w). However, that method also supports the use of a single contrast, training in a similar manner to self-SR methods.
@@ 3 Methods
Our proposed ECLARE improves resolution along the through-plane axis of an anisotropic 2D-acquired multi-slice MR image volume. To explain ECLARE, we first model the relationship between HR and LR volumes (Sec. 3.1). We then describe how to approximate the slice selection profile (Sec. 3.2) and FOV-aware resampling (Sec. 3.3). We approximate the slice selection profile with ESPRESO [14] and use it to train a CNN with in-plane HR patches and their simulated LR counterparts (Sec. 3.4). Finally, we show how ECLARE uses this 2D CNN to super-resolve the 3D volume (Sec. 3.5) and describe implementation details (Sec. 3.6).
@@ 3.2 Approximating the imaging process
3.3 FOV-aware resampling
While resampling is routine in digital signal processing, the exact implementation has several freedoms. ECLARE requires both downsampling (for training data simulation) and upsampling (for inference) and we would like both to be stable with respect to the FOV. We consider three conditions that are desirable to meet with respect to the FOV of the signal:
@@ 3.3 FOV-aware resampling
Resampling algorithms cannot meet all of these conditions in general since only a discrete number of samples is permitted. However, since it is possible to satisfy two of these conditions, a compromise on the third is necessary. We quickly review popular modern implementations. The method scipy.ndimage.zoom [49] meets conditions 1) and 3). It adjusts the sample spacing to preserve the FOV center while locking the locations of boundary samples. The method torch.nn.functional.interpolate [36] meets only condition 2). It preserves the sample spacing, but neither the FOV center nor the boundary sample locations. Instead, it anchors the extent on the first sample. We place more importance on the center of the FOV and the sample spacing and therefore desire conditions 1) and 2). We summarize the resampling method proposed in [15] that maintains the FOV center and sample spacing while sacrificing the precise location of the boundary samples. A comparison of the three implementations is shown in Fig. 2.
Table 1: Quantitative results for the ablation study. For brevity, E denotes ESPRESO, F denotes FOV-aware resampling, and W denotes the modified WDSR architecture. The model absent of any contributions is Baseline, and the model using all contributions is ECLARE. See Sec. 4.2 for complete details. PSNR and SSIM are reported as mean $\pm$ standard deviation. Best results are bolded and second-best results are underlined.
Configuration E F W PSNR SSIM
Baseline $27.72\pm 0.84$ $0.84\pm 0.02$
$28.00\pm 0.86$ $0.85\pm 0.02$
$27.91\pm 0.86$ $0.85\pm 0.02$
$27.48\pm 2.80$ $0.80\pm 0.16$
$29.28\pm 0.85$ $0.87\pm 0.02$
$29.32\pm 0.85$ $0.86\pm 0.02$
$28.61\pm 0.87$ $0.87\pm 0.02$
ECLARE $\mathbf{30.29\pm 0.87}$ $\mathbf{0.88\pm 0.02}$
Figure 4: Ablation study results labeled by configuration. E denotes ESPRESO, F denotes FOV-aware resampling, and W denotes the modified WDSR architecture. A sagittal slice of a representative subject is shown, and a zoomed inset indicates regions where the absence of ESPRESO in the algorithm yields a hyperintense artifact.
@@
As in [15], we introduce a fixed coordinate system for the signal before and after resampling. Let the position of the first pixel be centered at zero, let $N$ be the number of pixels before resampling, and let the spacing before resampling be one. The width of each pixel is equal to the spacing; before resampling it is therefore also one. The first pixel $p_{0}$ is centered at zero and the last pixel $p_{N-1}$ is centered at N1N-1. Since each pixel has a width of one, $p_{0}$ extends from 0.5-0.5 to $0.5$ and $p_{N-1}$ extends from N1.5N-1.5 to N0.5N-0.5. Therefore the entire image extends from 0.5-0.5 to N0.5N-0.5.
We now calculate the new positions and extents of pixels after resampling with respect to the same coordinate system. Let the new spacing (and pixel width) be $d$. After resampling, the new image must have an integer number of pixels. With $N^{\prime}$ as the number of pixels after resampling and $\lfloor\cdot\rceil$ as the rounding operator,
@@
Accordingly, we notate the first pixel after resampling as $p^{\prime}_{0}$ and the last pixel as $p^{\prime}_{N^{\prime}-1}$.
In order to center around the same FOV, the center position of the entire image should be identical before and after resampling. Since the center of the FOV is the halfway point between the first and last pixels, we can write
@@
Finally, we can calculate the new extents of the entire image after resampling. The first pixel extends from $p^{\prime}_{0}-\frac{d}{2}$ to $p^{\prime}_{0}+\frac{d}{2}$ and the last pixel extends from $p^{\prime}_{N-1}-\frac{d}{2}$ to $p^{\prime}_{N-1}+\frac{d}{2}$. Thus, the image extents are $p^{\prime}_{0}-\frac{d}{2}$ to $p^{\prime}_{N-1}+\frac{d}{2}$.
The result of this method yields an FOV-aware resampling implementation that maintains the FOV center and guarantees the desired pixel spacing. This is true in both upsampling and downsampling scenarios as demonstrated in Fig. 2.
@@ 3.4 Training a deep learning model with paired data
We now develop a relationship between in-plane and through-plane data and train a SR network. Paired LR and HR data are necessary for supervised learning. We extract HR 2D patches from in-plane and produce blurred 2D patches using Eq. 1 with the learned $\tilde{h}$ from ESPRESO. These blurred 2D patches are then downsampled using the FOV-aware resampling described in Sec. 3.3. Thus, each LR patch has the appropriate simulated slice thickness, separation, and FOV.
@@
where $\lceil\cdot\rceil$ is the ceiling operation. Accordingly, to achieve the desired $r$, we first upsample by $\lceil r\rceil$ and then downsample by $\frac{r}{\lceil r\rceil}$. Upsampling is done with pixel shuffle and downsampling is done with FOV-aware resampling (Sec. 3.3). Our model is illustrated in Fig. 3.
@@ 4.1 Data and evaluation metrics
Experiments using pairs of 2D-acquired LR images and 3D-acquired HR images would be ideal. However, quantitative analysis is not feasible without simulations. Acquiring high-resolution 2D scans increases scan time linearly unless methods to reduce scan time are used. Longer scan times and acceleration techniques can introduce artifacts that would be detrimental to the analysis. 3D scans are quicker, but acquire data differently and sometimes require different contrasts, eliminating possible comparisons. This motivates a physics-based simulation approach, established as a precedent previously [37, 41, 19, 61, 63, 62, 13, 14].
Our experiments began with HR isotropic 3D-acquired image volumes of resolution $1\times 1\times 1\text{~mm}^{3}$. These images were then degraded with Eq. 1, setting $h$ to the shape of the slice excitation profile created by the Shinnar-Le Roux (SLR) pulse design algorithm [37, 41], $s$ to the desired slice spacing (described below), and $x$ to a 3D-acquired isotropic volume. The implementation of the SLR algorithm we use is from the sigpy library [33].
The physical dimensions of the slice thickness is determined by the FWHM of the slice selection profile. Since our images began with $1\text{~mm}$ thick slices, this involved choosing the FWHM of $h$ to be equal to the desired slice thickness. In clinical neuroimaging, it is common to acquire no slice gap or a slice gap equal to $1\text{~mm}$ or $1/3$ the slice thickness, so we chose to simulate 2D acquisitions of thicknesses $3,4,\text{ and }5\text{~mm}$ with gaps $0,1,1.2,\text{ and }1.5\text{~mm}$, yielding a total of $8$ resolutions. We adopt the notation $T\|G$ [39] for a thickness of $T$ mm and gap of $G$ mm and list our choices here: $(3\|0),(3\|1),(4\|0),(4\|1),(4\|1.2),(5\|0),(5\|1),\text{ and }(5\|1.5)$. These simulation experiments were conducted on two datasets. The first consisted of $50$ T1-w images from healthy subjects in OASIS-3 [23]. The second consisted of $30$ T2-w FLuid Attenuated Inversion Recovery (FLAIR) images from people with MS in the 3D-MR-MS dataset [24].
@@
4.2 Ablation on components
In this section, we perform an ablation to validate components of our method. The ablation cohort consisted of $50$ subjects from the OASIS-3 dataset at the resolution $4\|1.2$. Since we have three primary technical contributions (FOV-aware resampling, the use of ESPRESO to estimate slice selection profiles, and modifications to the WDSR architecture for non-integer scale factors), there are eight configurations to examine. When FOV-aware resampling is absent, torch.interpolate with cubic B-splines is used instead. When ESPRESO is absent, a Gaussian slice selection profile with thickness equal to the slice separation is used instead. When the WDSR architectural modifications are absent, resampling is used instead of pixel shuffle. (This choice of resampling is determined by the corresponding ablation and refers to either FOV-aware resampling or torch.interpolate.) The Baseline configuration omits all of these changes, while the ECLARE uses all contributions.
We evaluated each configuration with PSNR and SSIM, shown in Table 1. By both metrics, ECLARE outperformed all other configurations. With the Wilcoxon signed-rank test and Bonferroni correction for multiple comparisons, the improved performance of ECLARE over all other ablations was significant ($p\ll 0.001$). However, the ablation study revealed some interesting contributions. First, we will discuss the single-contribution configurations. In all three cases, results did not substantially improve from the baseline for PSNR and SSIM. Interestingly, using ESPRESO alone without FOV-aware resampling or the WDSR pixel shuffle modifications reduced quantitative performance; note how the PSNR score exhibits not only lower average PSNR but higher PSNR standard deviation as well. This potentially could signify the sensitivity of ESPRESO to the resampling procedure, which is indicated by the two-contribution ablations. For the two-contribution ablations, performance improved in all cases compared to the Baseline. The combination of ESPRESO with either of the other two contributions yielded the highest PSNR results, and the scores for SSIM are comparable across all three two-contribution ablations.
Qualitative results of a representative subject for the ablation study are shown in Fig. 4. Both the baseline and using ESPRESO alone failed to recover from the blocky aliasing artifacts. The inclusion of ESPRESO is important, however. Note the regions indicated by the zoomed insets; when ESPRESO is absent from a configuration, hyperintense artifacts appear at the interfaces between dark and light intensities. However, when ESPRESO is present, those hyperintense artifacts are not apparent.
@@ 5 Discussion and conclusions
In this paper we presented ECLARE, a novel self-SR algorithm for anisotropic volumes that trains solely on data extracted from in-plane slices and thus requires no external training data. We presented a model of the imaging process in 2D multi-slice MRI and used ESPRESO [14] to estimate the slice selection profile. With an estimated slice profile, slice thickness and slice separation are both modeled, allowing for training data simulation from the HR in-plane data already present in the anisotropic volume. ECLARE uses a neural network architecture based on WDSR to efficiently train weights from scratch in a short amount of time, taking approximately 5 minutes on a single NVIDIA V100 GPU with a memory footprint under 2 GB. The result is a SR technique that respects the LR image contrast and presence of pathology, does not require external training data, and produces HR image volumes with the correct digital resolution and FOV. In the ablation study, we observed that the individual technical contributions of ECLARE are not independently additive, but instead exhibit strong interdependencies. In particular, ESPRESO alone did not improve quantitative performance and, when used without appropriate resampling, could even degrade PSNR and SSIM. However, when combined with either FOV-aware resampling or the modified WDSR architecture, ESPRESO consistently led to improved performance, suggesting that accurate modeling of slice selection profiles is most effective when paired with resampling strategies that respect acquisition geometry. These results indicate that the full ECLARE framework benefits from the complementary interaction of all three components, rather than from any single contribution in isolation. When compared to four other approaches on two simulated datasets, our approach achieved superior quantitative and qualitative results with respect to signal recovery metrics PSNR and SSIM as well as a downstream task metric—cDSC for brain region and MS lesion segmentation. This is true for all investigated resolutions except for $3\|0$, where SMORE performs best. This is likely due to two reasons: 1) SMORE does not have a model of slice gap, and therefore will perform best when no gap is present in the data; and 2) SMORE relies on pre-trained weights and was primarily tested and evaluated using data with resolutions $2\|0$ and $3\|0$.
We note that computation time is a primary limitation of ECLARE. ECLARE must fit a model for each multi-slice 2D MR volume. Although its training time is relatively fast compared to other self-SR methods, it is slower than pre-trained models that perform inference only. It would be possible to pre-train-then-fine-tune ECLARE as SMORE, but this introduces a tradeoff between time and relying on pre-training as a prior more than the in-plane data itself. Future work will investigate these tradeoffs in detail. Additionally, future work will involve experimental design to evaluate the viability of super-resolution in clinical practice.
@@
Code, Data, and Materials Availability
The code associated with this manuscript is open-source and available at
https://www.github.com/sremedios/eclare. The code associated with the comparison methods SMORE [62], single-contrast INR [34], and SynthSR [18], as well as the code for the two downstream brain parcellation methods SLANT [17] and SynthSeg [2] are available from their respective online sources. The data used in this study, OASIS-3 [23] and 3D-MR-MS [24], are open-source and available from their respective online sources.